Database last updated on Sun, 01 Feb 2026 19:40:06 +0000 by localhost
Instrument constant changes for the week before 2023235 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 28 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
aode
ds
ed
fz
hg
sl
zs
235
CHURCHILL
#011
menu,Next schedule will be ecds (now calds ZS= 93 ZS= 93.268)
74
74
2
2
11
2
5
(0 sec)
Dead time (0 sec)
Nothing to report.
Standard lamp R6 (1 sec)
File name
Count
Serial
Expected (+/-20)
Measured
30-day history
B23523.011
2
#011
2091
2130